# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v1

@@ III-B1 Semantic Analysis and Injection Point Identification
The process begins with semantic and structural analysis of the benign code. An LLM-based agent analyzes the function’s control and data flow to infer its purpose, input/output behavior, and internal logic (Code Understanding agent). It then qualitatively identifies potential sources—locations where untrusted input may enter the program—and sinks—locations where sensitive operations occur (Sanitization Detector agent). The agent also detects sanitization logic that would prevent the intended vulnerability from manifesting, such as bounds checks or input validation.
This process mirrors traditional symbolic static analysis and taint-tracking techniques [36], but applied inversely: rather than proving the presence of a vulnerability, we infer the barriers that prevent it, and identify how they could be weakened or removed. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
III-B2 RAG-Based Vulnerability Injection with Fine-Tuning
@@
In the following, we first explain the retrieval-augmented vulnerability generation, then detail the low-rank adaptation fine-tuning method, followed by supervised fine-tuning and reinforcement learning strategies.
@@
Once candidate injection points are identified, an agent transforms the code to inject the target vulnerability (Vulnerability Injector agent). This agent operates under a retrieval-augmented generation (RAG) setup. Given an input function $c^{b}$, the agent retrieves a set of similar benign/vulnerable pairs $\{c^{b_{(j)}},c^{v_{(j)}}\}_{j=1}^{k}$ from a knowledge base using embedding-based similarity. Retrieved examples are aligned using line-level diffs, marking the injected edits.
@@
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [31]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size.
Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$.
This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
@@
where each benign function $c^{b_{(j)}}$ is accompanied by a structured prompt $q^{(j)}$. The prompt is generated by prior workflow steps from III-B1, incorporating (1) semantic and structural analysis results of the benign code, (2) the original benign function itself, and (3) relevant examples from the Retrieval-Augmented Generation (RAG) knowledge base, annotated to highlight vulnerability injection locations. The objective is for the agent to generate the corresponding vulnerable function $\hat{c}^{v}_{j}$ given this prompt.
@@
III-C Validation Module
The validation module ensures that injected code $c^{v}$ meaningfully introduces the intended vulnerability into the original benign function $c^{b}$. This process involves both deterministic tools and language model agents, and follows the workflow depicted in Figure 1.
First, a diff agent computes a token-level difference between $c^{v}$ and $c^{b}$ (Difference agent). If the difference is trivial (e.g., limited to whitespace, comments, or reformatting), the workflow returns to the injection stage with corrective feedback (Vuln. Injector Identical Code agent).
If the change is substantive, a vulnerability verification agent (LLM-based) assesses whether the injected code actually reflects the vulnerability that was intended (Diff. Analyzer agent). This agent is provided with rich context, including the original function $c^{b}$, the modified function $c^{v}$, their annotated difference, the vulnerability class to inject, and a self-rationale generated by the injection agent, which explains what edits were made and why they introduce the vulnerability. This enables the agent to evaluate not just the presence of code edits, but also whether those edits semantically fulfill the injection objective.
The modified code is analyzed using cppcheck [^1] https://cppcheck.sourceforge.io/. A validation agent then interprets its diagnostic report alongside prior context to assess whether the injected vulnerability aligns with the intended CWE (Vuln. Verifier).
If validation fails at this stage, the code is passed to a refinement agent, which revises the output based on prior feedback. The updated code is re-evaluated through the same pipeline. To ensure termination, the workflow is bounded by a fixed number of correction iterations: 10.
@@ IV Experimental Setup
  • RQ2: What impact does fine-tuning the vulnerability injection agent have on AVIATOR’s performance?
@@ IV-A Benchmark Construction
  • SARD test suite 100 [16, 40]: Contains a set of 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [33]. Due to the compact size and self-contained structure of these examples, we use SARD 100 to perform rapid fully automated validation.
  • FormAI [32]: Includes more realistic and structurally complex synthetic examples. While it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for manual validation, offering a stronger challenge to our injection framework.
  • PrimeVul [11]: Contains over 5,000 real-world C/C++ code pairs annotated with vulnerability labels. Due to the lack of full context (e.g., missing class definitions, global variables, and interprocedural dependencies), PrimeVul is not suitable for automated formal verification. Instead, we use it for (1) fine-tuning the injection model, (2) constructing the retrieval-augmented generation (RAG) knowledge base, and (3) manually evaluating injection success on 45 randomly sampled pairs.
@@ IV-B Implementation Details
We use gte-Qwen2-1.5B-Instruct for embedding vulnerable and benign function pairs in the RAG knowledge base. This embedder achieves top scores on the MTEB benchmark for embeddings of up to 8K tokens [43]. Since 99% of our training dataset from PrimeVul fits within this limit, we avoid chunking. Retrieval parameters were set at the recommended $k=4$.
@@
We fine-tune the vulnerability injection agent using 3.5k aligned pairs from the PrimeVul training set. For supervised fine-tuning (SFT), we trained for 5 epochs (loss convergence) with a learning rate of $2\times 10^{-4}$, batch size of 4, and LoRA ranks varying between 16 and 256. For reinforcement learning with GRPO, we trained the model for 1 epoch (reward convergence), using a learning rate of $2\times 10^{-4}$, batch size of 8, and generating four completions per policy iteration. The KL regularization coefficient $\beta$ (0.04) and clipping parameter $\epsilon$ (0.2) followed the recommended default values from GRPO implementation using Adam optimizer.
@@
AVIATOR is implemented using Python as well as HuggingFace’s Transformers and Parameter-Efficient Fine-Tuning libraries. All experiments were run on NVIDIA A100 or H100 80GB GPUs.
IV-C Ablation Study Design
We conduct an ablation study to quantify the contribution of each agentic component in the AVIATOR workflow. To assess the necessity and impact of this modular structure, we construct and evaluate six progressively simplified variants by incrementally removing components from the full workflow.
@@ IV-C Ablation Study Design
  • 5-Agent (Injection-Only Module), $W_{5}$: Implements the full injection module (Code Understanding, Information Loader, Sanitization Detector, Vulnerability Injector, Polisher) without validation (Figure 1, top).
@@
  • 9-Agent, $W_{9}$: Adds the cppcheck static analysis tool and corresponding Verifier Agent to further validate the generated vulnerability and refine injection quality based on external tool feedback.
@@ IV-D Evaluation and Metrics
For PrimeVul, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [21, 10, 29, 46]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) undecidable—typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.
@@ V Results and Analysis
FormAI Automated $AISR_{5}$ 89%
@@
Answer to RQ2. Fine-tuning the vulnerability injection agent using SFT leads to consistent and significant performance gains for the whole AVIATOR workflow, particularly on complex datasets like FormAI. GRPO fails to match this improvement despite being substantially more resource-intensive. These results highlight the efficiency and generalization strength of SFT, even when trained on an entirely separate dataset.
@@
To evaluate the contribution of AVIATOR’s agentic workflow design, we performed an ablation study on seven progressively extended configurations ($W_{1}$$W_{13}$), incrementally adding specialized agents for analysis, validation, and refinement. We additionally compare four model variants: (i) base Qwen2.5-Coder, (ii) Qwen2.5 fine-tuned using SFT, (iii) Qwen2.5 fine-tuned using GRPO, and (iv) the general-purpose Llama-4-Maverick. Figure 2 shows the average injection success rate across 5 trials ($AISR_{5}$) on both SARD100 and FormAI.
@@
Refer to caption Refer to caption
(a) SARD 100 (b) FormAI
Figure 2: AVIATOR average injection success rate on 5 attempts. Ablation study for LLM agents using Llama-4-Maverick, Qwen2.5-Coder without fine-tuning and Qwen2.5-Coder with the vulnerability injection agent fine-tuned using SFT and GRPO.
@@ VI Discussion
This section reflects on why AVIATOR works, its practical impact, limitations, and directions for future research.
@@ VI-A Why Does AVIATOR Work?
(1) AVIATOR benefits from agentic decomposition. The injection task is divided among 13 specialized agents, each handling specific operations like vulnerability localization, injection, sanitization detection, or diff-based verification. This modular approach enhances accuracy significantly, as demonstrated by improvements from 31% to 91% on FormAI and from 84% to 94% on SARD100.
@@
(3) AVIATOR leverages iterative workflow refinement to improve injection robustness. Agents operate in feedback loops to enhance injection robustness. For example, the Diff Agent triggers refinement cycles when necessary, boosting accuracy up to 10 points on FormAI and 4 points on SARD100.
(4) AVIATOR incorporates traditional static analysis tools to validate injection quality. The Verifier Agent uses cppcheck to detect whether the injected code violates real-world safety constraints. Its feedback is interpreted and used by subsequent agents to iteratively refine the injection.

### v2

@@ III-C1 Code Understanding (LLM)
The process begins with semantic and structural analysis of the benign code. The agent analyzes the control and data flow of the input source code $c^{b}$ to infer its functional intent, input/output behavior, and internal logic. It also qualitatively identifies potential sources—locations where untrusted input may enter the program—and sinks—locations where sensitive operations occur. This stage mirrors the comprehension step a human analyst performs.
III-C2 Information Loader (tool)
Retrieves a detailed definition and description of the type of vulnerability that needs to be injected. This helps ensure the LLM receives proper context on this vulnerability category.
III-C3 Vuln.-Specific Code Understanding (LLM)
Examines the benign code to identify elements that prevent the intended vulnerability from manifesting. These include explicit sanitization checks (e.g., bounds validation) as well as broader safety mechanisms such as correct memory management, error handling, or defensive programming patterns. The agent produces a structured inventory of these safeguards, which downstream agents use to determine how the code can be realistically transformed into a vulnerable variant. This process mirrors traditional symbolic static analysis and taint-tracking techniques [37], but applied inversely. Instead of proving correctness, it highlights the logic that enforces safety. This reverse reasoning is critical to selecting feasible and non-trivial injection points.
III-C4 Vulnerability Injector (LLM + RAG + fine-tuning).
@@ III-C2 Information Loader (tool)
Relies on previous information to apply the transformation and justify its injection. To further boost the performance of this critical component we leverage retrieval-augmented generation (RAG) as well as fine-tuning. In the following, we first explain the retrieval-augmented vulnerability generation, then detail the low-rank adaptation fine-tuning method, followed by supervised fine-tuning and reinforcement learning strategies.
@@
Once candidate injection points are identified, an agent transforms the code to inject the target vulnerability (Vulnerability Injector agent). This agent operates under a retrieval-augmented generation (RAG) setup. Given an input function $c^{b}$, the agent retrieves a set of $k$ similar benign/vulnerable pairs $\{c^{b_{(j)}},c^{v_{(j)}}\}_{j=1}^{k}$ from a knowledge base using embedding-based similarity. Retrieved examples are aligned using line-level diffs, marking the injected edits.
@@
To specialize the vulnerability injection agent without incurring the computational cost of full model fine-tuning, we apply Low-Rank Adaptation (LoRA) [33]. This technique is particularly effective in settings where the base model is large, but the task-specific dataset is limited in size. Rather than updating the full weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA learns a low-rank decomposition by injecting two trainable matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$, into the model’s architecture: $W=W_{0}+BA$. This reduces the number of trainable parameters while retaining the benefits of large-scale pretraining.
@@
where each benign function $c^{b_{(j)}}$ is accompanied by a structured prompt $q^{(j)}$. The prompt is generated by using the output of prior workflow steps III-C1, III-C2, III-C3. This prompt incorporates (1) semantic and structural analysis results of the benign code, (2) the original benign function itself, and (3) relevant examples from the Retrieval-Augmented Generation (RAG) knowledge base, annotated to highlight vulnerability injection locations. The objective is for the agent to generate the corresponding vulnerable function $\hat{c}^{v}_{j}$ given this prompt.
@@
III-C5 Polisher (LLM)
After a first version of the injection is completed, this agent was introduced to fix common issues from the Vulnerability Injector. It helps to normalize formatting, corrects truncated outputs, and removes superficial edits such as comment-only changes. The polished candidate is then handed to the Validation Module.
III-D Validation Module
The Validation Module ensures that a candidate $c^{v}$ produced by the Injection Module meaningfully introduces the intended vulnerability into the original function $c^{b}$. It mirrors the review stage of human analysts: checking plausibility, and revising when necessary. The module combines difference analysis, contextual LLM judgment, static analysis, and bounded correction loops (Figure 1).
III-D1 Difference (tool): Triviality Gate
Given $(c^{b},c^{v})$, compute a line diff. If there are no changes or the changes are only whitespace/comments, the injection failed. In that case, the workflow is routed to the Vuln. Injector Identical Code. Otherwise, if there is at least one non-trivial edit, the candidate proceeds directly to the Diff Refiner.
III-D2 Vuln. Injector Identical Code (LLM)
Invoked only when the triviality gate detects no or trivial edits only. This agent re-attempt to generate $c^{v}$ from scratch, incorporating the context provided by prior agents, with the aim of increasing the likelihood that the agent actually introduces the vulnerability this time. The new candidate is returned to the Difference (Triviality Gate) for re-evaluation. This loop is bounded (default: 10 attempts) to prevent cycling.
III-D3 Diff Refiner (LLM)
For candidates with substantive edits, the Diff Refiner analyzes the generated code together with the annotated difference and determines whether the changes correctly implement the requested vulnerability. If the edits are incomplete, incorrect, or misaligned, the agent updates the vulnerable code accordingly. Its judgment draws on the context accumulated from all previous agents, ensuring that the revised $c^{v}$ both preserves original functionality and aligns with the CWE intent before proceeding to downstream validation.
III-D4 Difference (tool): Reviewer Diff
The revised $c^{v}$ is diffed again to produce the reviewer-oriented, annotated difference consumed by following analyzers.
III-D5 Diff Analyzer (LLM)
The Diff Analyzer serves as a second reviewer to double-check the edits introduced in $c^{v}$. It analyzes the generated code together with the annotated difference and reassesses whether the modifications correctly introduce the requested vulnerability while preserving the original functionality.
III-D6 Lightweight Static Analyzer (tool)
Runs lightweight static analysis on $c^{v}$ using cppcheck [^1] https://cppcheck.sourceforge.io/. Diagnostics are passed to the verifier. While we use cppcheck for C/C++, this component is generic and can be swapped with a static analyzer appropriate for other languages.
III-D7 Vulnerability Verifier (LLM)
Integrates diagnostics from the static analyzer with the annotated diff and CWE context to decide whether the candidate indeed reflects the intended flaw. On success, the module outputs the validated $c^{v}$; otherwise, control transfers to the fixer.
III-D8 Vulnerability Fixer (LLM) and Controlled Restart
Attempts a constrained repair using all accumulated context (code, diffs, CWE, rationales, diagnostics). The repaired candidate re-enters validation at the Lightweight Static Analyzer agent up to one time. If the intended vulnerability still does not materialize, the controller restarts from the initial Vulnerability Injector, exploiting LLM stochasticity to explore a different edit trajectory. To ensure termination, the workflow is bounded by a fixed number of correction iterations: 10.
@@ IV Experimental Setup
  • RQ2: What is the contribution of fine-tuning the vulnerability injection agent to AVIATOR’s overall performance?
@@ IV-A Benchmark Construction
  • SARD test suite 100 [17, 41]: Contains 102 compact, synthetic test cases, including 34 vulnerable/benign function pairs that can be detected through formal verification tools such as ESBMC [35]. Their compact and self-contained design, makes them ideal for rapid, fully automated validation and served as our primary development benchmark.
  • FormAI [34]: Includes more realistic and structurally complex synthetic examples. Since it does not provide aligned benign/vulnerable pairs, we manually created benign versions for 37 randomly sampled vulnerable functions, without altering the program’s intent and ensuring no vulnerabilities were detectable by formal verification tools on the resulting benign versions. ESBMC is again used for manual validation, offering a stronger challenge to our injection framework.
  • PrimeVul [11]: Contains over 5,000 real-world C/C++ code pairs annotated with vulnerability labels. Due to the lack of full context (e.g., missing class definitions, global variables, and interprocedural dependencies), as well as bigger code samples, formal verification tools such as ESBMC are not directly applicable. We use the provided: (1) training split of 3,789 pairs across 112 CWEs to fine-tune the Vulnerability Injector, (2) validation split of 480 pairs (63 CWEs) as the retrieval knowledge base in AVIATOR’s RAG setup, and (3) 45 randomly sampled pairs from the test split (24 CWEs) for manual evaluation of injection success.
@@ IV-B Implementation Details
We use gte-Qwen2-1.5B-Instruct as the embedding model for vulnerable/benign code function pairs drawn from the validation split of PrimeVul (§IV-A) that constitutes the RAG knowledge base. At the time of our experiments, this model achieved state-of-the-art scores on the MTEB benchmark for embeddings up to 8K tokens [44]. Since 99% of functions in the PrimeVul training split fall below this length, no chunking was required. We set the retrieval parameter to $k=4$, following the recommended default.
The Vulnerability Description used by the Information Loader agent (§III-C2) covers 133 CWE categories and explains common programming mistakes that lead to each vulnerability type.
@@
We fine-tune the vulnerability injection agent using 3,789 aligned pairs across 112 CWEs from the PrimeVul training set. For supervised fine-tuning (SFT), we trained for 5 epochs (loss convergence) with a learning rate of $2\times 10^{-4}$, batch size of 4, and LoRA ranks varying between 16 and 256. For reinforcement learning with GRPO, we trained the model for 1 epoch (reward convergence), using a learning rate of $2\times 10^{-4}$, batch size of 8, and generating 4 completions per policy iteration. The KL regularization coefficient $\beta$ (0.04) and clipping parameter $\epsilon$ (0.2) followed the recommended default values from GRPO implementation using Adam optimizer.
@@
AVIATOR is implemented using Python as well as HuggingFace’s Transformers and Parameter-Efficient Fine-Tuning libraries, and VLLM for inference. All experiments were run on NVIDIA A100 or H100 80GB GPUs.
IV-C Empirical Study Design
We conduct an empirical study to quantify the contribution of adding agentic components to the AVIATOR workflow. The goal is to assess both the necessity and the impact of the modular structure by systematically simplifying the framework and comparing outcomes. To this end, we construct and evaluate six progressively reduced variants of AVIATOR, obtained by incrementally removing agents from the full pipeline.
@@ IV-C Empirical Study Design
  • 5-Agent (Injection-Only Module), $W_{5}$: Implements the full injection module (Code Understanding, Information Loader, Vuln.-Specific Code Understanding, Vulnerability Injector, Polisher) without validation (Figure 1, top).
@@
  • 9-Agent, $W_{9}$: Adds the static analysis tool and corresponding Verifier Agent to further validate the generated vulnerability and refine injection quality based on external tool feedback.
@@ IV-D Evaluation and Metrics
For PrimeVul, where formal verification tools such as ESBMC are not directly applicable due to the lack of full context (e.g., missing class definitions and global variables), as well as bigger code samples, we manually validated a sample of 45 injected functions drawn uniformly at random, following prior work [23, 10, 31, 47]. The evaluation was performed by three authors experienced in C/C++, including two senior security researchers. Each injected sample was categorized as (1) weakness present, (2) no weakness, or (3) undecidable—typically due to reliance on external functions or context beyond the injected function. Since we lack complete test cases for most programs, we assess the presence of weaknesses rather than confirmed vulnerabilities (which require demonstrable exploitability). The reported injection success rate reflects the percentage of samples found to contain a weakness, based only on those that could be conclusively assessed.
@@ V Results and Analysis
FormAI Automated $AISR_{5}$ 91%
@@
It is important to note that, even without fine-tuning, AVIATOR achieves strong baseline performance. This indicates that the agentic workflow itself, through decomposition, correction loops, and validation, is sufficient to generate realistic and diverse vulnerabilities. Fine-tuning provide further gains, but is not strictly necessary for the framework to operate effectively. While baseline performance depends on how well the underlying LLM has been trained on the target language, this property remains especially valuable when extending to programming languages or vulnerability classes where aligned training data is scarce, since AVIATOR can still produce rich vulnerability patterns.
@@
Answer to RQ2. Fine-tuning the vulnerability injection agent using SFT leads to consistent and significant performance gains for the whole AVIATOR workflow, particularly on complex datasets like FormAI. GRPO fails to match this improvement despite being substantially more resource-intensive. These results highlight the efficiency and generalization strength of SFT, even when trained on an entirely separate dataset. Additionally, the results show strong baseline performance without fine-tuning, indicating that the agentic workflow alone can inject vulnerabilities. This is essential for transfer to new programming languages or CWE categories where aligned data is unavailable.
@@
To evaluate the contribution of AVIATOR’s agentic workflow design, we conducted an empirical study on seven progressively extended configurations ($W_{1}$$W_{13}$), incrementally adding specialized agents for analysis, validation, and refinement. We additionally compare four model variants: (i) base Qwen2.5-Coder, (ii) Qwen2.5 fine-tuned using SFT, (iii) Qwen2.5 fine-tuned using GRPO, and (iv) the general-purpose Llama-4-Maverick. Figure 2 shows the average injection success rate across 5 trials ($AISR_{5}$) on both SARD100 and FormAI.
Refer to caption Refer to caption
(a) SARD 100 (b) FormAI
Figure 2: AVIATOR average injection success rate on 5 attempts. Empirical study for LLM agents using Llama-4-Maverick, Qwen2.5-Coder without fine-tuning and Qwen2.5-Coder with the vulnerability injection agent fine-tuned using SFT and GRPO.
@@
V-D Compute and Cost Estimates
On 8 NVIDIA A100 GPUs, running 128 jobs in parallel, we generated 3,200 vulnerable samples in 2h57m. This averages to 3s per injection with concurrency. At Maryland’s commercial electricity rate (14.82¢/kWh), this corresponds to a cost of about \$0.44 per 1,000 injections (GPU power only). These numbers show that AVIATOR can effectively scale dataset generation at very low cost.
@@ VI Discussion
This section reflects on why AVIATOR works, its practical impact, limitations, and outlook.
@@ VI-A Why Does AVIATOR Work?
(1) AVIATOR benefits from agentic decomposition. The injection task is divided among 13 specialized agents, each handling specific operations like vulnerability localization, injection, or diff-based verification. This modular approach enhances accuracy significantly, as demonstrated by improvements from 31% to 91% on FormAI and from 84% to 95% on SARD100.
@@
(3) AVIATOR leverages iterative workflow refinement to improve injection robustness. Agents operate in feedback loops to enhance injection robustness. For example, the Diff Agent triggers refinement cycles when necessary, boosting accuracy up to 10 points on FormAI and 4 points on SARD100. Notably, AVIATOR achieves a strong baseline even without fine-tuning, provided the underlying LLM has sufficient coverage of the target language, highlighting that its workflow design is the primary driver of performance. This makes the framework readily portable to new programming languages and under-studied vulnerability classes, while ensuring diversity in the types and structures of injected flaws.
(4) AVIATOR integrates traditional static analysis tools to validate injection quality. The Verifier Agent uses cppcheck to detect whether the injected code violates real-world safety constraints. Its feedback is interpreted and used by subsequent agents to iteratively refine the injection.